A scratcher has four even numbers and four odd numbers covered. What are the odds that you would scratch off three even numbers on your first try?
Answer by stanbon(75887)

A scratcher has four even numbers and four odd numbers covered. What are the odds that you would scratch off three even numbers on your first try?
-----------
P(3 even with no replacement) = (4/8)(3/7)(2/6) = 0.0714...
P(not 3 even) = 0.92857..
----------
Odds in favor of 3 even = P(3 even)/P(not 3 even) = 0.07689...
===========================
Cheers,
Stan H.